# Pre-Chirp-Domain Index Modulation for Full-Diversity Affine Frequency Division Multiplexing towards 6G

### v3

@@ IV-B Diversity Analysis
$\Pr\big{(}[\mathbf{x},\mathbf{\Phi}]\to[\hat{\mathbf{x}},\hat{\mathbf{\Phi}}]\big{)}$
@@ V Parameter Optimization
In this section, we establish an optimization problem for the optimal $c_{2}$ alphabet design, where the PSO algorithm is employed, to enhance the BER performance.
@@ V-A Problem Formulation
$\hskip 17.07182pt\quad c_{2}\in\mathbb{I},$ (48d)
where the objective function can be expressed as
$O_{k,j}(\mathcal{P}_{c})\!=\!\sum_{\mathbf{x}^{\prime}\!,\mathbf{x}}\sum_{r=1}^{\mathcal{R}}\sum_{p=1}^{P}\!\!\sum_{n=0}^{N-1}\!\left(1\!-\!\cos\left(\psi_{n}+\theta_{n}^{\prime}\!-\!\theta_{n}\right)\right),$ (49)
in which $\psi_{n}$ represents the phase difference between $x_{\operatorname{loc}_{p}+n}^{\prime}$ and $\big{(}x_{\operatorname{loc}_{p}+n}\big{)}^{*}$, i.e.,
$e^{\textsf{j}\psi_{n}}=x_{\operatorname{loc}_{p}+n}^{\prime}\big{(}x_{\operatorname{loc}_{p}+n}\big{)}^{*}$ (50)
with $x_{\mathrm{loc}_{p}+n}^{\prime}$ and $x_{\mathrm{loc}_{p}+n}$ denoting the elements of $\mathbf{x}^{\prime}$ and $\mathbf{x}$, respectively, and
@@
The derivation of $O_{k,j}(\mathcal{P}_{c})$ (49) is given in Appendix -B.
@@ V-B1 Transformation of $\mathbf{x}^{\prime}$ and $\mathbf{x}$
Since $x_{\operatorname{loc}_{p}+n}^{\prime}$ and $\big{(}x_{\operatorname{loc}_{p}+n}\big{)}^{*}$ take values from the $M$-PSK constellation, $\psi_{n}$ is given by
$\psi_{n}=\frac{2\pi k}{M},k=-(M-1),\ldots,-1,0,1,\ldots,M-1,$ (52)
where each possible value is assigned with equal probability, i.e., $\psi_{n}$ follows a discrete uniform distribution.
For the case where $\mathbf{x}^{\prime}\neq\mathbf{x}$, $O_{k,j}(\mathcal{P}_{c})$ can be expressed as
$O_{k,j}^{(\mathbf{x}^{\prime}\neq\mathbf{x})}(\mathcal{P}_{c})=\sum_{\mathbf{x}^{\prime}\neq\mathbf{x}}\sum_{r=1}^{\mathcal{R}}\sum_{p=1}^{P}\sum_{n=0}^{N-1}\left(-\cos\left(\psi_{n}+\theta_{n}^{\prime}-\theta_{n}\right)\right)$
@@
Therefore, for the original problem (48), it is sufficient to consider the case of $\mathbf{x}^{\prime}\!=\!\mathbf{x}$. In this case, $\psi_{n}\!=\!0$ for all $n\!=\!0,1,\ldots,N$, and the problem (48) can be expressed as
@@ V-C Problem Solver
5:Initialize local optimal position of each particle $\bm{p}_{n_{p},local}=\bm{p}_{n_{p}}^{(0)}$, calculate global optimal position $\bm{p}_{global}=\arg\max\limits_{1\leq n_{p}\leq N_{p}}\big{(}F_{p}(\bm{p}_{n_{p}}^{(0)})\big{)}$;
@@ VI Simulation Results and Analysis
In this section, we carry out simulations to evaluate the performance of the proposed AFDM-PIM schemes, and conduct the BER performance comparison between the proposed AFDM-PIM scheme and other existing counterparts. The accuracy of the PEP-based theoretical derivation is also investigated in comparison with the Monte Carlo simulation results. In simulations, the carrier frequency is set to $f_{c}\!=\!8$ GHz, and the subsymbol spacing in the DAF domain is set to $f_{s}\!=\!1.5$ kHz. Two distinct simulation scenarios are conducted, namely, when the full diversity condition (45) is satisfied and when it is not satisfied. The maximum normalized Doppler shifts for these two scenarios are set to $\alpha_{\mathrm{max}}\!=\!1$ and $\alpha_{\mathrm{max}}\!=\!2$, corresponding to the high-speed scenarios with the maximum speed of mobile station $v_{e}\!=\!202.5$ km/h and the ultra high-speed scenarios with the maximum speed of mobile station $v_{e}\!=\!405$ km/h, respectively. The Doppler shift of each path is generated according to Jakes Doppler spectrum approach as $\alpha_{p}\!=\!\alpha_{\mathrm{max}}\cos(\theta_{p,d})$, where $\theta_{p,d}\!\in\![-\pi,\,\pi]$ (for integer Doppler cases, the Doppler shift is $\lfloor\alpha_{\mathrm{max}}\cos(\theta_{p,d})\rfloor$). We set the pertinent parameters of Algorithm 1 as ${v}_{\max}\!=\!0.05$, $\varpi\!=\!0.5$, $N_{p}\!=\!200$, $I_{\mathrm{max}}\!=\!300$, $\varrho_{global}\!=\!2$ and $\varrho_{local}\!=\!2$. Unless otherwise specified, the ML detector is employed for both the proposed AFDM-PIM and the classical benchmarks.
@@ VI-A Effectiveness of Algorithm 1
First, the effectiveness of the proposed $c_{2}$ alphabet design with Algorithm 1 is validated. Fig. 4 compares the BER performance achieved by the $c_{2}$ alphabet designed by Algorithm 1 and that attained by the initial solution, namely, the first initial particle $\bm{p}_{1}^{(0)}$ for Algorithm 1, given two numbers of paths $P\!=\!2$ and 3. Binary PSK (BPSK) is employed as the modulation scheme for the data bits, and the parameter settings are $(N,G,\lambda,d_{\mathrm{max}},\alpha_{\mathrm{max}})\!=\!(6,2,3,1,1)$. Note that the full diversity order condition is satisfied, and consequently the BER performance improves with the increase in the number of paths. Fig. 4 indicates that the $c_{2}$ alphabet designed by Algorithm 1 exhibits superior performance by about 3 dB in SNR than the initial solution at the BER level of $10^{-3}$, given $P\!=\!3$. The results of Fig. 4 demonstrate the effectiveness of Algorithm 1.
Figure 5: Performance comparisons between the proposed AFDM-PIM scheme and the classical AFDM scheme given two different spectral efficiencies.
@@ VI-B Performance Comparison with Existing Benchmarks
Fig. 5 compares the BER performance of the proposed AFDM-PIM with that of the classical AFDM scheme [17] under the same doubly dispersive channel with $P\!=\!4$. The proposed AFDM-PIM employs BPSK and quadrature PSK (QPSK) to achieve the spectral efficiency of $2$ and $3$ bit/s/Hz, respectively. The other parameters are as $(N,G,\lambda,d_{\mathrm{max}},\alpha_{\mathrm{max}})\!=\!(8,2,4,2,2)$. To reach the same spectral efficiency levels in the classical AFDM, the number of subcarriers is set to $8$, and QPSK and $8$-PSK are utilized correspondingly. The both schemes do not satisfy the full diversity order condition. Specifically, the proposed AFDM-PIM does not satisfy Condition 1 in Theorem 3, while the AFDM does not satisfy Theorem 1 of [17]. It can be seen from Fig. 5 that the AFDM-PIM demonstrates an about $2$ dB gain in the SNR compared to the AFDM scheme at the BER level of $10^{-3}$. This is because the AFDM-PIM allows lower order constellations to reach the same spectral efficiency as the AFDM. The results of Fig. 5 indicates that the AFDM-PIM offers a viable alternative for communication under doubly dispersive channels.
Figure 6: Performance comparisons between the proposed AFDM-PIM and the AFDM-IM scheme given two different spectral efficiencies.
In Fig. 6, we evaluate the BER performance of the proposed AFDM-PIM and the AFDM-IM [29] under a doubly dispersive channel with the number of paths $P\!=\!3$ and given two spectral efficiencies of $2$ and $3$ bit/s/Hz. For the proposed AFDM-PIM, $(N,G,\lambda,d_{\mathrm{max}},\alpha_{\mathrm{max}})$ are set to $(8,2,4,1,2)$, and BPSK and QPSK are employed to reach the spectral efficiencies of $2$ and $3$ bit/s/Hz, respectively. For the AFDM-IM scheme, each group comprises $n$ subcarriers, with $a$ subcarriers activated at each transmission. To reach the same spectral efficiency levels for the AFDM-IM, $(n,a)\!=\!(4,2)$ and $(n,a)\!=\!(8,7)$ are employed, respectively, and $8$-PSK is utilized. Note that the both schemes do not meet the full diversity order condition. It can be seen from Fig. 6 that the AFDM-PIM scheme yields better BER performance than the AFDM-IM scheme, with an SNR gain of about $2$ dB at the BER level of $10^{-3}$. The enhanced performance is attributable to the capability of the AFDM-PIM to employ lower order constellations than the AFDM-PIM, while reaching the same spectral efficiency as the AFDM-IM.

### v4

@@ IV-B Diversity Analysis
$\Pr\big{(}[\mathbf{x},\mathbf{\Phi}]\to[\hat{\mathbf{x}},\hat{\mathbf{\Phi}}]\big{)}\approx\!\frac{1}{12}\prod_{\iota=1}^{\kappa}\frac{4PN_{0}}{\lambda_{\iota}^{(\mathbf{\Psi})}}\!+\!\frac{1}{4}\prod_{\iota=1}^{\kappa}\frac{3PN_{0}}{\lambda_{\iota}^{(\mathbf{\Psi})}}$
@@
IV-C Spectral Efficiency Analysis
This subsection examines the SE of the proposed AFDM-PIM scheme. Classic AFDM and IM-assisted AFDM (referred to as AFDM-IM) are utilized as benchmarks to highlight the superiority of the proposed approach.
The SEs of the classic AFDM [21] and AFDM-IM [40] are given by
$S_{\mathrm{AFDM}}=\log_{2}(M)$ (46)
and
$S_{\mathrm{AFDM}-\mathrm{IM}}=\frac{\left\lfloor\log_{2}\binom{N_{c}}{a}\right\rfloor+a\log_{2}(M)}{N_{c}},$ (47)
where $N_{c}$ and $a$ represent the numbers of subcarriers in each group and activated subcarriers at each transmission, respectively, for the AFDM-IM scheme.
Based on the description in Subsection III-B, the SE of the proposed AFDM-PIM can be calculated as
$S_{\mathrm{AFDM}-\mathrm{PIM}}=\frac{\left\lfloor\log_{2}(N_{c}!)\right\rfloor}{N_{c}}+\log_{2}(M).$ (48)
As can be seen from the above analysis, the proposed AFDM-PIM demonstrates a higher SE than AFDM, due to the presence of index modulation. Furthermore, in contrast to AFDM-IM, which employs inactive subcarriers at each transmission, AFDM-PIM exhibits a higher SE, attributed to the capacity of all subcarriers to carry modulated data. In summary, compared with classic methods, the proposed AFDM-PIM scheme offers an enhancement in terms of SE, without incurring any additional hardware cost.
@@ V Parameter Optimization
In this section, we establish an optimization problem for the optimal $c_{2}$ alphabet design, where the PSO algorithm is employed to enhance the BER performance.
@@ V-A Problem Formulation
$\hskip 17.07182pt\quad c_{2}\in\mathbb{I}.$ (51d)
To derive the objective function, consider the elements of $\mathbf{x}^{\prime}$ and $\mathbf{x}$, $x_{\mathrm{loc}_{p}+n}^{\prime}$ and $x_{\mathrm{loc}_{p}+n}$, which can be expressed as $x_{\mathrm{loc}_{p}+n}^{\prime}\!=\!r_{p,n}e^{\textsf{j}\vartheta_{p,n}}$ and $x_{\mathrm{loc}_{p}+n}\!=\!r_{p,n}^{\prime}e^{\textsf{j}\vartheta_{p,n}^{\prime}}$. Define
@@
and
[…]
where $\psi_{p,n}\!=\!\vartheta_{p,n}^{\prime}\!-\!\vartheta_{p,n}$. Then the objective function $O_{k,j}(\mathcal{P}_{c})$ can be expressed as
$O_{k,j}(\mathcal{P}_{c})\!=\!\sum_{\mathbf{x}^{\prime}\!,\mathbf{x}}\sum_{r=1}^{\mathcal{R}}\sum_{p=1}^{P}\!\sum_{n=0}^{N-1}F_{p,n}.$ (54)
The derivation of (54) is given in Appendix -B.
@@ V-B1 Transformation of $\mathbf{x}^{\prime}$ and $\mathbf{x}$
Consider the constant modulus constellation using phase shift keying (PSK) as an example. Since $x_{\operatorname{loc}_{p}+n}^{\prime}$ and $x_{\operatorname{loc}_{p}+n}$ take values from the $M$-PSK constellation, $r_{p,n}\!=\!r_{p,n}^{\prime}\!=1$, and $\psi_{p,n}$ is given by
$\psi_{p,n}=\frac{2\pi k}{M},k=-(M-1),\ldots,-1,0,1,\ldots,M-1,$ (55)
where each possible value is assigned with equal probability, i.e., $\psi_{p,n}$ follows a discrete uniform distribution. For the case where $\mathbf{x}^{\prime}\!\neq\!\mathbf{x}$, $O_{k,j}(\mathcal{P}_{c})$ can be expressed as
$O_{k,j}^{(\mathbf{x}^{\prime}\neq\mathbf{x})}(\mathcal{P}_{c})=\sum_{\mathbf{x}^{\prime}\neq\mathbf{x}}\sum_{r=1}^{\mathcal{R}}\sum_{p=1}^{P}\sum_{n=0}^{N-1}\left(-\cos\left(\psi_{p,n}+\theta_{n}^{\prime}-\theta_{n}\right)\right)$
@@
Therefore, for the objective function $O_{k,j}(\mathcal{P}_{c})$, it is sufficient to consider the case of $\mathbf{x}^{\prime}\!=\!\mathbf{x}$. In this case, $\psi_{p,n}\!=\!0$ for all $n\!=\!0,1,\ldots,N-1$, and $O_{k,j}(\mathcal{P}_{c})$ can be expressed as
$O_{k,j}(\mathcal{P}_{c})=\sum_{r=1}^{\mathcal{R}}\sum_{p=1}^{P}\sum_{n=0}^{N-1}\left(1-\cos\left(\theta_{n}^{\prime}-\theta_{n}\right)\right).$ (57)
For non-constant modulus constellations, take quadrature amplitude modulation (QAM) as an example. Then $x_{\operatorname{loc}_{p}+n}^{\prime}$ and $x_{\operatorname{loc}_{p}+n}$ take values from the $M$-QAM constellation and the constellation diagram is symmetric with respect to the origin. Hence, for each pair of $x_{\operatorname{loc}_{p}+n}^{\prime}$ and $x_{\operatorname{loc}_{p}+n}$ with phase difference $\psi_{p,n}=\gamma$, there exist a pair of corresponding constellation points with phase difference $\psi_{\hat{p},\hat{n}}=\gamma+\pi$. Therefore, for the case of $\mathbf{x}^{\prime}\neq\mathbf{x}$, we can obtain
$\sum_{\mathbf{x}^{\prime}\neq\mathbf{x}}\sum_{r=1}^{\mathcal{R}}\sum_{p=1}^{P}\sum_{n=0}^{N-1}\left(r_{p,n}r_{p,n}^{\prime}\cos\left(\theta_{n}^{\prime}\!-\!\theta_{n}+\psi_{p,n}\right)\right)$
$=\sum_{r=1}^{\mathcal{R}}\sum_{p=1}^{P}\sum_{n=0}^{N-1}\!\Big{(}\!\sum_{\psi_{p,n}}\!\!\cos\left(\gamma\!+\!\theta_{n}^{\prime}\!-\!\theta_{n}\right)\!+\!\cos\left(\gamma\!+\!\pi\!+\!\theta_{n}^{\prime}\!-\!\theta_{n}\right)\!\Big{)}$
$=0.$ (58)
Consequently, it is sufficient to focus on the case of $\mathbf{x}^{\prime}\!=\!\mathbf{x}$ for the objective function $O_{k,j}(\mathcal{P}_{c})$. Under this condition, $r_{p,n}\!=\!r_{p,n}^{\prime}$ and $\psi_{p,n}\!=\!0$, and $O_{k,j}(\mathcal{P}_{c})$ can be expressed as
$O_{k,j}(\mathcal{P}_{c})\!=\!\sum_{\mathbf{x}^{\prime},\mathbf{x}}\sum_{r=1}^{\mathcal{R}}\sum_{p=1}^{P}\sum_{n=0}^{N-1}2r_{p,n}^{2}\left(1-\cos\left(\theta_{n}^{\prime}-\theta_{n}\right)\right),$ (59)
where signal amplitudes $r_{p,n}$ are unrelated to optimization.
In summary, for both constant modulus and non-constant modulus constellations, the optimization problem (51) can be transformed into
@@ V-C Problem Solver
5:Initialize local optimal position of each particle $\bm{p}_{n_{p},local}=\bm{p}_{n_{p}}^{(0)}$, calculate global optimal position $\bm{p}_{global}=\arg\max\limits_{1\leq n_{p}\leq N_{p}}\big{(}{\cal F}_{p}(\bm{p}_{n_{p}}^{(0)})\big{)}$;
@@ VI Simulation Results and Analysis
In this section, we carry out simulations to evaluate the performance of the proposed AFDM-PIM schemes, and conduct the BER performance comparison between the proposed AFDM-PIM scheme and other existing counterparts. The accuracy of the PEP-based theoretical derivation is also investigated in comparison with the Monte Carlo simulation results. In simulations, the carrier frequency is set to $f_{c}\!=\!8$ GHz, and the subcarrier spacing in the DAF domain is set to $f_{s}\!=\!1.5$ kHz. Two distinct simulation scenarios are conducted, namely, when the full diversity condition (45) is satisfied and when it is not satisfied. The maximum normalized Doppler shifts for these two scenarios are set to $\alpha_{\mathrm{max}}\!=\!1$ and $\alpha_{\mathrm{max}}\!=\!2$, corresponding to the high-speed scenarios with the maximum speed of mobile station $v_{e}\!=\!202.5$ km/h and the ultra high-speed scenarios with the maximum speed of mobile station $v_{e}\!=\!405$ km/h, respectively. [^1] The relationship between speed and normalized Doppler shift is shown in Appendix -C. The Doppler shift of each path is generated according to Jakes Doppler spectrum approach as $\alpha_{p}\!=\!\alpha_{\mathrm{max}}\cos(\theta_{p,d})$, where $\theta_{p,d}\!\in\![-\pi,\,\pi]$ (for integer Doppler cases, the Doppler shift is $\lfloor\alpha_{\mathrm{max}}\cos(\theta_{p,d})\rfloor$). We set the pertinent parameters of Algorithm 1 as ${v}_{\max}\!=\!0.05$, $\varpi\!=\!0.5$, $N_{p}\!=\!200$, $I_{\mathrm{max}}\!=\!300$, $\varrho_{global}\!=\!2$ and $\varrho_{local}\!=\!2$. The adopted simulation parameters are listed in Table III. Unless otherwise specified, the ML detector is employed for both the proposed AFDM-PIM and the classical benchmarks.
TABLE III: Simulation Parameters
Parameter Description Value
$c$ Speed of light in free space $3\times 10^{8}$ m/s
$f_{c}$ Carrier frequency 8 GHz
$f_{s}$ Subcarrier spacing 1.5 kHz
$v_{e}$ Maximum speed of mobile station 202.5/405 km/h
${v}_{\max}$ Maximum particle velocity 0.05
$\mathcal{P}_{c}$ Alphabet of $c_{2}$ realizations for $\lambda$ = 2/3/4 {0.20, 0.60} {0.29, 0.62, 0.93} {0.01, 0.20, 0.41, 0.80}
$\varpi$ Inertia weight 0.5
$N_{p}$ Number of particles 200
$I_{\max}$ Maximum iteration value 300
$\varrho_{global}$ Global updating coefficients 2
$\varrho_{local}$ Local updating coefficients 2
@@ VI-A Effectiveness of Algorithm 1
We first validate the effectiveness of the proposed $c_{2}$ alphabet design with Algorithm 1. Fig. 4 compares the BER performance achieved by the $c_{2}$ alphabet designed by Algorithm 1 and that attained by the initial solution, namely, the first initial particle $\bm{p}_{1}^{(0)}$ for Algorithm 1, given two numbers of paths $P\!=\!2$ and 3. Binary PSK (BPSK) is employed as the modulation scheme for the data bits, and the parameter settings are $(N,G,\lambda,d_{\mathrm{max}},\alpha_{\mathrm{max}})\!=\!(6,2,3,1,1)$. Since the full diversity order condition is satisfied, the BER performance improves with the increase in the number of paths. Fig. 4 indicates that the $c_{2}$ alphabet designed by Algorithm 1 exhibits superior performance by about 3 dB in SNR than the initial solution at the BER level of $10^{-3}$, given $P\!=\!3$. The results of Fig. 4 demonstrate the effectiveness of Algorithm 1.
Figure 5: Performance comparison between the proposed AFDM-PIM scheme and the OFDM-IM scheme given two different spectral efficiencies.
Figure 6: Performance comparison between the proposed AFDM-PIM scheme and the classical AFDM scheme given two different spectral efficiencies.
@@ VI-B Performance Comparison with Existing Benchmarks
Fig. 5 compares the BER performance achieved by the proposed AFDM-PIM and the OFDM-IM [33] under a doubly dispersive channel with $P\!=\!4$ and given two SEs of $2$ and $3$ bit/s/Hz. For the proposed AFDM-PIM, $(N,G,\lambda,d_{\mathrm{max}},\alpha_{\mathrm{max}})$ are set to $(8,2,4,0,1)$, and BPSK and quadrature PSK (QPSK) are employed to reach the SEs of $2$ and $3$ bit/s/Hz, respectively. For the OFDM-IM scheme [33], each group comprises $n$ subcarriers, with $a$ subcarriers activated at each transmission. To reach the SEs of $2$ and $3$ bit/s/Hz, $(n,a)\!=\!(4,2)$ and $(n,a)\!=\!(8,7)$ are employed, respectively, and $8$-PSK is utilized. It can be seen from Fig. 5 that at the BER level of $10^{-3}$, our AFDM-PIM attain the SNR gains of about 5 dB and 4 dB over the OFDM-IM for the cases of $\text{SE}\!=\!3$ bits/s/Hz and 2 bits/s/Hz, respectively. This enhanced performance is attributable to the intrinsic advantage of AFDM-PIM, which can achieve full diversity order on doubly dispersive channels.
Fig. 6 compares the BER performance of our AFDM-PIM with that of the classical AFDM scheme [21] under the same doubly dispersive channel with $P\!=\!4$. The proposed AFDM-PIM employs BPSK and QPSK to achieve the SEs of $2$ and $3$ bit/s/Hz, respectively. The other parameters are as $(N,G,\lambda,d_{\mathrm{max}},\alpha_{\mathrm{max}})\!=\!(8,2,4,2,2)$. To reach the same SE levels in the classical AFDM [21], the number of subcarriers is set to $8$, and QPSK and $8$-PSK are utilized correspondingly. The both schemes do not satisfy the full diversity order condition. Specifically, the proposed AFDM-PIM does not satisfy Condition 1 in Theorem 3, while the AFDM does not satisfy Theorem 1 of [21]. It can be seen from Fig. 6 that the AFDM-PIM demonstrates an about $2$ dB gain in the SNR compared to the AFDM scheme at the BER level of $10^{-3}$. This is because the AFDM-PIM allows lower order constellations to reach the same SE as the AFDM. The results of Fig. 6 indicates that the AFDM-PIM offers a viable alternative for communication under doubly dispersive channels.
Figure 7: Performance comparison between the proposed AFDM-PIM scheme and the AFDM-IM scheme given two different spectral efficiencies.
In Fig. 7, we evaluate the BER performance of the proposed AFDM-PIM and the AFDM-IM [40] under a doubly dispersive channel with the number of paths $P\!=\!3$ and given two SEs of $2$ and $3$ bit/s/Hz. For the proposed AFDM-PIM, $(N,G,\lambda,d_{\mathrm{max}},\alpha_{\mathrm{max}})$ are set to $(8,2,4,1,2)$, and BPSK and QPSK are employed to reach the SEs of $2$ and $3$ bit/s/Hz, respectively. For the AFDM-IM scheme [40], each group comprises $n$ subcarriers, with $a$ subcarriers activated at each transmission. To reach the same SE levels for the AFDM-IM, $(n,a)\!=\!(4,2)$ and $(n,a)\!=\!(8,7)$ are employed, respectively, and $8$-PSK is utilized. Note that the both schemes do not meet the full diversity order condition. It can be seen from Fig. 7 that our AFDM-PIM scheme yields better BER performance than the AFDM-IM scheme, with an SNR gain of more than $2$ dB at the BER level of $10^{-3}$. The enhanced performance is attributable to the capability of our AFDM-PIM to employ lower order constellations than the AFDM-PIM, while reaching the same SE as the AFDM-IM.